ON A COMPUTER RECOGNITION OF 3-MANIFOLDS

SERGEI V. MATVEEV

Abstract. We describe theoretical backgrounds for a computer program that recognizes all closed orientable 3-manifolds up to complexity 8. The program can treat also not necessarily closed 3-manifolds of bigger complexities, but here some unrecognizable (by the program) 3-manifolds may occur.